jigglypuff314 (jigglypuff314):

Could someone help me with a Calc problem? :3 The position function for an object is given by s = 2t^3 - 3 (tan(pi/12 * t)) for t >_ 0 , where s is measured in feet and t is measured in seconds. At time t = 2, is the object speeding up or slowing down? Explain your reasoning.

jigglypuff314 (jigglypuff314):

\[s = 2t ^{3}-3\tan(\frac{ \pi }{ 12 } t ) ~~~~ for ~~t \ge 0\]

OpenStudy (anonymous):

ummmm

OpenStudy (agent0smith):

Differentiate the position function to find velocity, then again to find acceleration. Find the acceleration when t=2. Find the velocity when t=2. If a and v both have the same sign (ie both are positive or both negative), the object is speeding up. If they have different signs, it's slowing down.

jigglypuff314 (jigglypuff314):

okay, thank you! :)

jigglypuff314 (jigglypuff314):

@agent0smith for this problem I got that at t = 2 v = 22.953 and a = 23.683 so because a and v are both positive, the object is speeding up would that be correct?

OpenStudy (agent0smith):

Yep, if a and v have the same sign (ie same direction) the object is speeding up.

jigglypuff314 (jigglypuff314):

I was just not sure if I got the correct a and v values :/

OpenStudy (agent0smith):

You could check on wolfram alpha.

jigglypuff314 (jigglypuff314):

lol k, thanks